OpenStudy (anonymous):

If alfa,beta are the zeros of a polynomial such that alfa+beta=-6 and alfa*beta=-4,the write the polynomial ? i need the procedure.

OpenStudy (anonymous):

\[ x^2 - S x + P =0 \] Where S is the sum of the roots and P is the product.

OpenStudy (anonymous):

Ellas, I personally like the explanation of that: \[P(x)=(x-r_1)(x-r_2)=x^2-(r_1+r_2)x+r_1r_2\]

OpenStudy (anonymous):

\[ (x-\alpha)(x-\beta) =0\\ x^2 -(\alpha + \beta) x + \alpha \beta =0\\ x^2 + 6x -4=0 \]

OpenStudy (anonymous):

And, it is a trivial detail, but it is written 'alpha', not 'alfa'. (Though 'alfa' is more accurate with respect to how it is pronounced. The pronunciation of Greek letters and words is vastly different from that of the pronunciation of English letters and words.)

OpenStudy (anonymous):

yaya i got this

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

I do not get why you have to write a whole paragraph to make a distinction between alfa and alpha!!!

OpenStudy (anonymous):

Because it was fun to me.

OpenStudy (anonymous):

Have fun